OpenStudy (anonymous):

100 mL at 250°K. Which proportion could you use to find the volume v of the gas if the temperature decreases to 150°K?

OpenStudy (wolf1728):

According to Charles' Law V1 / T1 = V2 / T2 V2 = V1 * T2 / T1 V2 = 100 ml *150 / 250 V2 = 15,000 / 250 V2 = 60 ml And to recheck that: http://1728.org/charles.htm